1.m(7-2m)=6
2.1/x=x-2/24
3.4/x+10=4/x-1
4.x+3/x=x-4/3
5.40/x-20/x-3=8/7

Answer by stanbon(75887)

1.m(7-2m)=6
7m - 2m^2 = 6
2m^2-7m+6 = 0
2m^2-3m-4m+6 = 0
m(2m-3)-2(2m-3) = 0
(2m-3)(m-2) = 0
m = 3/2 or m=2
--------------------------
2. 1/x = (x-2)/24
24x = x-2
23x = -2
x = -2/23
--------------------------
3. 4/(x+10)=4/(x-1)
1/(x+10) = 1/(x-1)
x-1 = x+10
no solution
Your posting need parentheses.
-------------------
4.x+3/x=x-4/3
Needs parentheses.
5.40/x-20/x-3=8/7
need parentheses
----------------------
Cheers,
Stan H.